With an uncontrolled 1-phase half-wave rectifier with a phase voltage of 120 V.
A 10 Ω load is fed. Neglecting circuit losses,
a) Find the output voltage and current.
b) Find the effective phase current.
1
Expert's answer
2020-11-09T14:18:46-0500

Input voltage (rms value) "=120V"


Load resistance value "=10\varOmega"


a) Average voltage output is given


"V_o = \frac{V_m}{\pi}= \frac{120\times \sqrt{2}}{\pi} = 54.02V"


Average load current is given by


"I_o = \frac{V_o}{R} =\frac{54.02}{10}=5.4A"


b) Effective voltage or rms voltage output


"V_r = \frac{V_m}{2}=\frac{120\times\sqrt{2}}{2}=84.85V"


Effective output current


"I_r =\frac{V_r}{R}=\frac{84.85}{10}=8.5A"